INTRACLUSTER REACTIONS IN Ni-CO_{2} CLUSTERS

\underline{JONATHON A. MANER}, MATTHEW D. McDOWELL and MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, Georgia 30602..

Ni-CO_{2} cluster cations were produced by laser vaporization of a Ni rod in a pulsed CO_{2} expansion. Clusters were massselected in a time-of-flight mass spectrometer and studied with infrared laser photodissociation spectroscopy using an infrared OPO laser system. The infrared spectra of larger clusters provide evidence for an intracluster reaction assisted by CO_{2} solvation. These spectra are compared to determine the reaction product.